### v1

@@ 1 Introduction
This article proceeds as follows. In the next section, we review previous literature and delineate the contributions of our study. In Section 3, we provide a conceptual framework for our testing procedure that we formally propose in Section 4. Based on an extensive Monte Carlo study whose set-up is elaborated in Section 5, we examine the performance of our testing method and compare it with that of existing procedures in Section 6. After conducting another simulation study that offers additional intuition behind our testing method in Section 7, we conclude in Section 8.
@@ 2 Related literature
3 Firm conduct and the strength of BLP-style instruments
Setup: Discrete choice in differentiated goods markets
There are $J_{t}+1$ differentiated products in market $t=1,2,\ldots,T$ with size $M_{t}$, which measures the number of consumers. For product $j=0,1,\ldots,J_{t}$, there are observed attributes $(x_{jt},p_{jt})$ and an unobserved component $\xi_{jt}$. Observed attributes are grouped into two parts: (i) the exogenous part, $x_{jt}$, which is uncorrelated with unobservables $\bm{\xi}_{t}=(\xi_{0t},\xi_{1t},\ldots,\xi_{Jt})$ and (ii) the endogenous price, $p_{jt}$, which is correlated with $\bm{\xi}_{t}$. For each product, there are $R$ exogenous attributes in $x_{jt}=(x_{jt}^{(1)},\ldots,x_{jt}^{(r)},\ldots,x_{jt}^{(R)})$. A market $t$ is defined by a set of products offered, denoted by $\{0\}\cup\mathscr{J}_{t}$, a set of attributes of these products, denoted by $\chi_{t}=\{(x_{0t},p_{0t},\xi_{0t}),\ldots,(x_{Jt},p_{Jt},\xi_{Jt})\}$, and a set of characteristics of consumers, denoted by $V_{t}$.
A consumer $i$ with characteristics $v_{i}\in V_{t}$ has the indirect utility from consuming good $j$ given by $u_{ijt}=U(x_{jt},p_{jt},\xi_{jt},v_{i})$. A consumer purchases the product that maximizes his or her utility. The predicted market share of product $j$, denoted by $s_{jt}$, is the fraction of market consumers who prefer good $j$ over the other $J_{t}$ products:
@@ 3 Firm conduct and the strength of BLP-style instruments
Characteristics of a consumer, $v_{i}$, are decomposed into a common part shared across all individuals and heterogeneous components: $\nu_{i}=(\nu_{ip},\nu_{i1},\ldots,\nu_{iR})$, which are random utility components capturing heterogeneous valuations of $(p_{jt},x_{jt})$, and $\varepsilon_{ijt}$. The most commonly used utility specification in the literature is given by:
$u_{ijt}=\delta_{jt}+\nu_{ip}p_{jt}+\sum\limits_{r}\nu_{ir}x_{jt}^{(r)}+\varepsilon_{ijt},$
where $\delta_{jt}=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ is the mean utility of the product, which also absorbs the common part of $v_{i}$. Since we do not directly observe heterogeneous components $\nu_{i}$ and $\varepsilon_{ijt}$ in aggregated market-level data, we assume their distributions. Let $\varepsilon_{ijt}$ be i.i.d. (across both consumers and products) extreme value distributed, while $\nu_{i}$ follows a mean-zero multivariate normal distribution. [^3] The demographic-specific valuations of $(x_{jt},p_{jt})$ can also be modelled by either drawing from an empirical distribution or imposing distributional assumptions on the demographics of market $t$.
We normalize the indirect utility from consuming product $j=0$ by setting $u_{ijt}=\varepsilon_{ijt}$ across all consumers and refer to this product as the outside option. Assuming that the two random components, $\varepsilon$ and $\nu$, are independent, the predicted market share of product $j$ is given by
@@
Berry94 demonstrated that, for a given non-linear parameter vector $\theta_{2}$, there exists a unique vector of mean utilities $\bm{\delta}_{t}$ that equates the observed market shares with the predicted market shares. Specifically, let $\mathbf{S}_{t}=(S_{0t},S_{1t},\ldots,S_{Jt})$ denote the vector of observed market shares of products in market $t$. Then, $\mathbf{s}_{t}(\bm{\delta}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=\mathbf{S}_{t}$. This relationship allows us to invert the demand system in market $t$ and solve for $\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ for all $j\in\mathscr{J}_{t}$. [^4] This invertibility result can also be obtained for a more general specification of $U(\mathbf{x}_{t},\mathbf{p}_{t},\bm{\xi}_{t},v_{i})$. Specifically, when the utility specification induces connected substitutes relationships among products in the market, invertibility holds. See berry2013connected for further details. Finally, for a given parameter vector $\theta=(\theta_{1},\theta_{2})$, where $\theta_{1}=(\alpha,\beta)$ represents the linear parameters, we compute $\xi_{jt}(\theta)=\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})-(x_{jt}\beta-\alpha p_{jt})$.
@@
$E[\xi_{jt}(\theta_{0})\mathbf{Z}_{jt}]=0,$
where $\mathbf{Z}_{jt}=(x_{jt},\mathbf{z}_{jt})$ is a vector of instruments, and $\mathbf{z}_{jt}$ is a vector of excluded instruments. For the price of product $p_{jt}$, Berry95 used the sum of attributes across own-firm products and the sum of attributes across other-firm products as excluded instruments. [^5] Importantly, this type of instrument is also crucial in addressing endogeneity arising from observed market shares and prices of other products. To see this point, note that $\xi_{jt}(\theta)$ is a function of $\mathbf{S}_{t}$ and $\mathbf{p}_{t}$. Hence, estimating the inverted demand system with a flexible utility specification requires instruments that are exogenous to these variables as well. See berry2014identification for identification results and instrument requirements in nonparametric settings. Specifically, $\mathbf{z}_{jt}=(\mathbf{z}^{own}_{jt},\mathbf{z}^{other}_{jt})$ is given by
@@
where $\mathscr{J}_{ft}\setminus\{j\}$ denotes the set of products, excluding $j$, offered by firm $f$ in market $t$. These instruments, constructed as functions of product attributes for all products in the market, are referred to as BLP-style instruments throughout this article.
Firm conduct and BLP-style instruments
There are $F_{t}$ firms in market $t$. Importantly, firms may engage in collusion, internalizing the profits of products from their colluding partners, while the true industry conduct is unknown to the researcher or regulatory authority. The variable profit of firm $f=1,2,\ldots,F_{t}$ is given by:
$\Pi_{ft}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t},\phi)$ $=\sum_{j\in\mathscr{J}_{ft}}(p_{j}-mc_{j})\cdot s_{jt}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t})\cdot M_{t}$ (3)
@@
where $f_{c}$ is a firm index that treats colluding entities as a single entity. For instance, if firms 1 and 2 form a collusion, then $f_{c}=1$ for $f=1$ and $f=2$ resulting in $f_{c}=1,3,\ldots,F_{t}$.
The degree of internalization is measured by $\phi\in[0,1]$. As shown in (3), firm $f$ partially internalizes the profits of products owned by its colluding partners, weighting them by $\phi$ as part of its own profits. Firms in market $t$ set profit-maximizing prices simultaneously, based on the variable profits defined in (3), where the latter term is unknown to the researcher and regulatory authority. The Bertrand-Nash equilibrium prices in market $t$, and consequently the vector of equilibrium markups, are determined by the following First Order Conditions (FOCs):
@@
where $\mathbf{mc}_{t}$ is a vector of marginal costs of products, $\odot$ is the component-wise multiplication, and $\mathscr{H}_{t}$ is an unknown (to the researcher) $J_{t}\times J_{t}$ ownership index. For example, $\mathscr{H}_{t,jk}=1$ if $j,k\in\mathscr{J}_{ft}$ and 00 otherwise ($\phi=0$) in the case of competition, while $\mathscr{H}_{t}$ is a matrix of ones ($\phi=1$) in the case of full collusion. This markup equation demonstrates that in an oligopoly, pricing is influenced by the proximity of a product to its substitutes within the product characteristics space. Products facing close competition tend to have lower markups and prices, while those that are significantly differentiated can command higher markups and prices. More importantly, it highlights that the impact of these characteristics on pricing depends on the true ownership index.
@@
To clarify this point, assume a simple logit model, where $v_{ip}=v_{ir}=0$ in (1). Under this assumption, the equilibrium markups for firm $f$ in market $t$ are derived as:
@@
$\mathbf{p}_{ft}-\mathbf{mc}_{ft}=\left\{\begin{array}[]{ll}\frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\sum_{k\in\text{own}}X_{kt}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under competition }\vspace{0.1in}\\ \frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\big{(}\sum_{k\in\text{own}}X_{kt}+\sum_{k\in\text{colluding}}X_{kt}\big{)}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under collusion}\\ \end{array}\right.$ (7)
@@
where $\mathbf{1}_{ft}$ is a $J_{ft}\times 1$ unit vector and $X_{kt}=\exp(\delta_{kt})$. Note that $\delta_{kt}$ is a function of attributes of products $k$. One can see that the BLP-style instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$ in (2) for example, affect the equilibrium markups and hence, the equilibrium prices. Moreover, since the equilibrium markup is a function of $\mathscr{H}_{t}$, the identifying power of BLP-style instruments relies not only on the true market structure and competition level, but also on whether the industry conduct assumption upon which BLP-style instruments are constructed is correct. [^6] It also depends on variation in product attributes across products and markets and their distributional characteristics (e.g., right-skewness and symmetry). If the assumption is incorrect, then these instruments fail to properly capture the markup variation and, as a result, would have weaker identifying power. That is, when firms collude, colluding firms’ product attributes should be included in $\mathbf{z}^{own}$ rather than in $\mathbf{z}^{other}$.
@@
As an illustration, consider an automobile market with two firms, Volkswagen and Toyota, and three SUVs: Atlas (A), Tiguan (T), and Highlander (H). Given that the first two SUVs are owned by Volkswagen and the third by Toyota, the ownership matrix under Bertrand-Nash competition, $\mathscr{H}^{B}$, and the ownership matrix under full collusion of the two firms, $\mathscr{H}^{C}$, are represented as follows:
@@
As derived in Appendix A, assuming that $\alpha=\beta=1$ and $X_{A}=X_{T}=X_{H}=0.2$, [^7] The market share of the outside option is 62.5%. Also, results are qualitatively robust to different values of $\alpha$, $\beta$, and $X$s. we obtain
@@
$\left(\begin{array}[]{ccc}\frac{\partial p_{A}}{\partial x_{A}}&\frac{\partial p_{A}}{\partial x_{T}}&\frac{\partial p_{A}}{\partial x_{H}}\\ \end{array}\right)=\left\{\begin{array}[]{ccrl}(0.124&0.124&-0.037)&\textrm{under competition}\vspace{0.1in}\\ (0.125&0.125&0.125)&\textrm{under collusion}\\ \end{array}\right.$
@@
Two points are noteworthy. First, under Bertrand-Nash competition, the price of Atlas is more significantly affected by a change in the attributes of Tiguan than by a change in those of Highlander, and the two price effects are opposite in direction (0.124 vs. -0.037). For instance, after improving the fuel efficiency of Tiguan, Volkswagen would increase the prices of both Atlas and Tiguan to capitalize on its strengthened competitive advantage. However, the firm would lower the price of Atlas in response to an improvement in the fuel efficiency of Toyota’s Highlander. Secondly, in the case of collusion, the price effect of Highlander’s fuel efficiency improvement is 0.125, which is more than three times larger in magnitude than its effect under price competition (-0.037) and aligns in direction with the effect on Atlas’s price due to a change in the fuel efficiency of Tiguan. As a matter of fact, the three price effects under collusion are the same in magnitude because the three SUVs are equally substitutable for one another in this example.
@@
These observations indicate that, in the case of price competition, the sum of own-firm product attributes (competition IVs) has stronger identifying power than the sum of both own- and other-firm product attributes (collusion IVs). In contrast, if the two firms collude, collusion IVs exhibit a stronger correlation with price than competition IVs. Therefore, comparing the strength of competition IVs and collusion IVs could indicate whether firms in a market are competitive or collusive.
4 Testing frameworks

### v2

@@ Abstract
This version: March 1st 2025
First version: January 8th 2025
@@ 1 Introduction
We apply our testing method to study firm conduct in two differentiated product markets in South Korea: the new passenger car market and the instant noodles market. Our test rejects the hypotheses of brand-level profit maximization and price collusion in favor of joint profit maximization at the parent company level in the car market. In addition, our results suggest that, despite suspicions from the Korean Fair Trade Commission, instant noodle manufacturers have not engaged in any price collusion. Our findings align with anecdotal evidence of firm conduct in each market, court rulings, and previous empirical findings that instant noodle prices remain significantly below the collusive level (kim2024collusion).
This article proceeds as follows. In the next section, we review previous literature and delineate the contributions of our study. In Section 3, we provide a conceptual framework for our testing procedure which we formally propose in Section 4. Based on an extensive Monte Carlo study whose setup is elaborated in Section 5, we examine the performance of our testing method and compare it with that of existing procedures in Section 6. We then conduct another simulation study in Section 7 to provide further intuition behind our approach. In Section 8, we apply our method to study industry conduct in two differentiated product markets. Finally, we conclude in Section 9.
@@ 2 Related literature
3 Conceptual framework
3.1 Discrete choice in differentiated product markets
There are $J_{t}+1$ differentiated products in market $t=1,2,\ldots,T$. For product $j=0,1,\ldots,J_{t}$, there are observed attributes $(x_{jt},p_{jt})$ and an unobserved component $\xi_{jt}$. The observed attributes are grouped into two parts: (i) the exogenous part, $x_{jt}$, which is uncorrelated with the unobservables $\bm{\xi}_{t}=(\xi_{0t},\xi_{1t},\ldots,\xi_{Jt})$ and (ii) the endogenous price, $p_{jt}$, which is correlated with $\bm{\xi}_{t}$. For each product, there are $R$ exogenous attributes in $x_{jt}=(x_{jt}^{(1)},\ldots,x_{jt}^{(r)},\ldots,x_{jt}^{(R)})$. Market $t$ is defined by the set of products offered, denoted by $\{0\}\cup\mathscr{J}_{t}$, the set of attributes of these products, denoted by $\chi_{t}=\{(x_{0t},p_{0t},\xi_{0t}),\ldots,(x_{Jt},p_{Jt},\xi_{Jt})\}$, and the set of characteristics of consumers, denoted by $V_{t}$.
A consumer $i$ with characteristics $v_{i}\in V_{t}$ has the indirect utility from consuming product $j$ given by $u_{ijt}=U(x_{jt},p_{jt},\xi_{jt},v_{i})$. A consumer purchases the product that maximizes his or her utility. The predicted market share of product $j$, denoted by $s_{jt}$, is the fraction of market consumers who prefer good $j$ over all other $J_{t}$ products:
@@ 3.1 Discrete choice in differentiated product markets
Consumer characteristics, $v_{i}$, are decomposed into a common component shared across all individuals and heterogeneous components: (i) random utility components, $\nu_{i}=(\nu_{ip},\nu_{i1},\ldots,\nu_{iR})$, capturing heterogeneous valuations of $(p_{jt},x_{jt})$; and (ii) an idiosyncratic taste shock, $\varepsilon_{ijt}$. The most commonly used utility specification in the literature is given by:
$u_{ijt}=\delta_{jt}+\nu_{ip}p_{jt}+\sum\limits_{r}\nu_{ir}x_{jt}^{(r)}+\varepsilon_{ijt},$ (1)
where $\delta_{jt}=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ is the mean utility of the product, which also absorbs the common component of $v_{i}$. Since the heterogeneous components $\nu_{i}$ and $\varepsilon_{ijt}$ are not directly observed in aggregated market-level data, we assume their distributions. Let $\nu_{i}$ follow a mean-zero multivariate normal distribution, [^3] Demographic-specific valuations of $(x_{jt},p_{jt})$ can also be modeled either by drawing from an empirical distribution or by imposing distributional assumptions on the demographics of market $t$. while $\varepsilon_{ijt}$ are i.i.d. (across both consumers and products) extreme value distributed.
We normalize the indirect utility from consuming product $j=0$ by setting $u_{i0t}=\varepsilon_{i0t}$ for all consumers and refer to it as the outside option. Assuming that the two random components, $\nu$ and $\varepsilon$, are independent, the predicted market share of product $j$ is given by
@@
Berry94 demonstrated that, for a given non-linear parameter vector $\theta_{2}$, there exists a unique vector of mean utilities $\bm{\delta}_{t}$ that equates the observed market shares with the predicted market shares. Specifically, let $\mathbf{S}_{t}=(S_{0t},S_{1t},\ldots,S_{Jt})$ denote the vector of observed market shares of products in market $t$. Then, $\mathbf{s}_{t}(\bm{\delta}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=\mathbf{S}_{t}$. This relationship allows us to invert the demand system in market $t$ and solve for $\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})=x_{jt}\beta-\alpha p_{jt}+\xi_{jt}$ for all $j\in\mathscr{J}_{t}$. [^4] This invertibility result also holds for more general specifications of $U(\mathbf{x}_{t},\mathbf{p}_{t},\bm{\xi}_{t},v_{i})$. Specifically, when the utility specification induces connected substitutes relationships among products in the market, invertibility is ensured. See berry2013connected for further details. Finally, for a given parameter vector $\theta=(\theta_{1},\theta_{2})$, where $\theta_{1}=(\alpha,\beta)$ represents the linear parameters, we compute $\xi_{jt}(\theta)=\delta_{jt}(\mathbf{S}_{t},\mathbf{p}_{t},\mathbf{x}_{t};\theta_{2})-(x_{jt}\beta-\alpha p_{jt})$.
@@
$E[\xi_{jt}(\theta_{0})|\mathbf{Z}_{jt}]=0,$
where $\mathbf{Z}_{jt}=(x_{jt},\mathbf{z}_{jt})$ is a vector of instruments, and $\mathbf{z}_{jt}$ is a vector of excluded instruments. For the price of product $p_{jt}$, Berry95 used the sum of attributes across own-firm products and the sum of attributes across other-firm products as excluded instruments. [^5] Importantly, these instruments also help address endogeneity arising from observed market shares and prices of other products. Specifically, $\xi_{jt}(\theta)$ is a function of $\mathbf{S}_{t}$ and $\mathbf{p}_{t}$. Therefore, estimating the inverted demand system with a flexible utility specification requires instruments that are exogenous to these variables as well. See berry2014identification for identification results and instrument requirements in nonparametric settings. Specifically, $\mathbf{z}_{jt}=(\mathbf{z}^{own}_{jt},\mathbf{z}^{other}_{jt})$ is given by
@@
where $\mathscr{J}_{ft}\setminus\{j\}$ denotes the set of products, excluding $j$, offered by firm $f$ in market $t$. These instruments, constructed as functions of all product attributes in the market, are referred to as BLP-style instruments throughout this article.
3.2 Oligopoly model
There are $F_{t}$ firms in market $t$, which has a size of $M_{t}$, measuring the number of consumers. Importantly, firms may engage in collusion, internalizing the profits of their colluding partners, while the true industry conduct is unknown to the researcher or regulatory authority. The variable profit of firm $f=1,2,\ldots,F_{t}$ is given by:
$\Pi_{ft}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t},\phi)$ $=\sum_{k\in\mathscr{J}_{ft}}(p_{k}-mc_{k})\cdot s_{kt}(\mathbf{p}_{t};\mathbf{x}_{t},\bm{\xi}_{t})\cdot M_{t}$ (4)
@@ 3.2 Oligopoly model
where $f_{c}$ is a firm index that treats colluding entities as a single entity. For instance, if firms 1 and 2 collude, then $f_{c}=1$ for $f=1$ and $f=2$, resulting in $f_{c}=1,3,\ldots,F_{t}$.
The degree of profit internalization is measured by $\phi\in[0,1]$. As shown in equation (4), firm $f$ partially internalizes the profits of its colluding partners, weighting them by $\phi$ as part of its own profit. Firms in market $t$ set profit-maximizing prices simultaneously. The Bertrand-Nash equilibrium prices in market $t$, and consequently, the vector of equilibrium markups, are determined by the following First-Order Conditions (FOCs):
@@
where $\mathbf{mc}_{t}$ is a vector of marginal costs, $\odot$ denotes component-wise multiplication, and $\mathscr{H}_{t}$ is an unknown (to the researcher) $J_{t}\times J_{t}$ ownership matrix. For example, $\mathscr{H}_{t,jk}=1$ if $j,k\in\mathscr{J}_{ft}$ and 00 otherwise in the case of competition ($\phi=0$), while $\mathscr{H}_{t}$ is a matrix of ones in the case of full collusion ($\phi=1$ and $f_{c}=1$ for all $f$). As shown in the markup equation (5) above, the endogeneity of prices arises from the fact that firms set these prices based on demand, making them functions of the unobservable components $\bm{\xi}$.
@@
The markup equation (5) also demonstrates that in an oligopoly, pricing is influenced by the proximity of a product to its substitutes within the product characteristics space. Products facing close competition tend to have lower markups and prices, while those that are significantly differentiated can command higher markups and prices. More importantly, the equation highlights that the impact of these characteristics on pricing depends on the unknown ownership structure.
3.3 Firm conduct and the strength of BLP-style instruments
Logit case
@@ 3.3 Firm conduct and the strength of BLP-style instruments
To clarify this point, assume a simple logit model where $v_{ip}=v_{ir}=0$ in the utility specification (1). Under this assumption, the equilibrium markups for firm $f$ in market $t$ are derived as:
@@
$\mathbf{p}_{ft}-\mathbf{mc}_{ft}=\left\{\begin{array}[]{ll}\frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\sum_{k\in\text{own}}X_{kt}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under competition}\vspace{0.1in}\\ \frac{1}{\alpha}\bigg{(}\frac{1+\sum_{k\in\text{market}}X_{kt}}{1+\sum_{k\in\text{market}}X_{kt}-\big{(}\sum_{k\in\text{own}}X_{kt}+\sum_{k\in\text{colluding}}X_{kt}\big{)}}\bigg{)}\cdot\mathbf{1}_{ft}&\textrm{under full collusion}\\ \end{array}\right.$ (8)
@@
where $\mathbf{1}_{ft}$ is a $J_{ft}\times 1$ unit vector and $X_{kt}=\exp(\delta_{kt})$. Note that $\delta_{kt}$ is a function of the attributes of product $k$. One can see that the BLP-style instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$, affect the equilibrium markups and consequently, the equilibrium prices. Moreover, since the equilibrium markup is a function of $\mathscr{H}_{t}$, the identifying power of BLP-style instruments depends not only on the true market structure and competition level, but also on whether the industry conduct assumption upon which the BLP-style instruments are constructed is correct. [^6] It also depends on variation in product attributes across products and markets as well as their distributional characteristics (e.g., right-skewness and symmetry). If the assumption is incorrect, these instruments fail to properly capture markup variation and, as a result, have weaker identifying power. That is, when firms collude, the attributes of colluding firms’ products should be included in $\mathbf{z}^{own}$ rather than in $\mathbf{z}^{other}$.
@@
As an illustration, consider an automobile market with two firms, Volkswagen and Toyota, and three SUVs: Atlas (A), Tiguan (T), and Highlander (H). Given that the first two SUVs are owned by Volkswagen and the third by Toyota, the ownership matrix under Bertrand-Nash competition, $\mathscr{H}^{B}$, and the ownership matrix under full collusion between the two firms, $\mathscr{H}^{C}$, are represented as follows:
@@
As derived in Appendix LABEL:app:_simple_logit, assuming that $\alpha=\beta=1$ and $X_{A}=X_{T}=X_{H}=0.2$, [^7] The market share of the outside option is 62.5%. Also, results are qualitatively robust to different values of $\alpha$, $\beta$, and $X$s. we obtain
@@
$\left(\begin{array}[]{ccc}\frac{\partial p_{A}}{\partial x_{A}}&\frac{\partial p_{A}}{\partial x_{T}}&\frac{\partial p_{A}}{\partial x_{H}}\\ \end{array}\right)=\left\{\begin{array}[]{ccrl}(0.124&0.124&-0.037)&\textrm{if~{}}\phi=0\vspace{0.1in}\\ (0.125&0.125&0.125)&\textrm{if ~{}}\phi=1\\ \end{array}\right.$
@@
Two points are noteworthy. First, under Bertrand-Nash competition, the price of the Atlas is more significantly affected by a change in the attributes of the Tiguan than by a change in those of the Highlander, and the two price effects are opposite in direction (0.124 vs -0.037). For instance, after improving the fuel efficiency of the Tiguan, Volkswagen would increase the prices of both the Atlas and the Tiguan to capitalize on its strengthened competitive advantage. However, the firm would lower the price of the Atlas in response to an improvement in the fuel efficiency of Toyota’s Highlander. Secondly, under full collusion, the price effect of an improvement in the Highlander’s fuel efficiency is $0.125$, which is more than three times larger in magnitude than its effect under price competition (-0.037) and aligns in direction with the effect on the Atlas’s price due to a change in the fuel efficiency of the Tiguan. In fact, the three price effects under collusion are identical in magnitude because the three SUVs are equally substitutable for one another in this example.
@@
These observations indicate that, under price competition, the sum of own-firm product attributes (competition IVs) has greater identifying power than the sum of both own- and other-firm product attributes (collusion IVs). In contrast, if the two firms collude, collusion IVs exhibit a stronger correlation with price than competition IVs. Therefore, comparing the strength of competition IVs and collusion IVs can provide insights into whether firms in a market are competitive or collusive.
Nested logit case
To assess whether the findings above hold in a more general setting, we assume that the idiosyncratic taste shock $\varepsilon_{ijt}$ in the indirect utility function (1) follows a nested logit structure, parameterized by the nesting parameter $\sigma\in[0,1)$. [^8] Specifically, $\varepsilon_{ijt}=\zeta_{igt}+(1-\sigma)\overline{\varepsilon}_{ijt}$, where $\overline{\varepsilon}_{ijt}$ follows an i.i.d. extreme value distribution and $\zeta_{igt}$ has a unique distribution such that $\varepsilon_{ijt}$ remains extreme value distributed (Cardell97). A higher value of $\sigma$ implies stronger substitution between products belonging to the same group, indexed by $g=0,1,\ldots,G$. The outside option ($j=0$) is the sole member of group $g=0$. We then simulate equilibrium outcomes for five markets, each consisting of six firms ($F=6$) and 60 inside products ($J=60$), categorized into three groups ($G=3$). In each market, the first three firms collude ($f_{c}=1$ for $f=1,2,3$) with varying degrees of profit internalization, $\phi\in\{0,0.2,0.5,0.8,1\}$; they compete in the first market and fully collude in the last market.
Our objective is to analyze the $J\times J$ Jacobian matrix, denoted by $\frac{\partial\mathbf{p}}{\partial\mathbf{x}}$, where the $(j,k)$ entry corresponds to $\frac{\partial p_{j}}{\partial x_{k}}$ evaluated at equilibrium. Specifically, we first examine whether, under competition ($\phi=0$), the following condition holds:
$\left[\frac{\partial\mathbf{p}}{\partial\mathbf{x}}\right]_{jk}\cdot\left[\frac{\partial\mathbf{p}}{\partial\mathbf{x}}\right]_{jk^{\prime}}<0\quad\text{for $k\in\mathscr{J}_{f(j)}$ and $k^{\prime}\notin\mathscr{J}_{f(j)}$,}$
where $f(j)$ indexes the firm offering product $j$. Next, we investigate how varying the degree of profit internalization, $\phi$, affects this relationship. Analytical and computational details for deriving the Jacobian matrix are provided in Appendix LABEL:app:_nested_logit.
Figure 1: Effects of own-firm and other-firm attributes on equilibrium prices
(a) Competition ($\phi=0$)
(b) Partial profit internalization ($\phi=0.2$)
(c) Partial profit internalization ($\phi=0.5$)
(d) Partial profit internalization ($\phi=0.8$)
(e) Full profit internalization ($\phi=1$)
Notes: The figure presents heatmaps of the $J\times J$ Jacobian matrices, $\frac{\partial\mathbf{p}}{\partial\mathbf{x}}$, derived from the nested logit specification under various values of the profit internalization parameter, $\phi$. See the main text for interpretation and Appendix LABEL:app:_nested_logit for details on the data-generating process used to compute equilibrium outcomes and Jacobian matrices.
Figure 1(e) presents heatmaps of the Jacobian matrix for the five markets. Each grid represents the $(j,k)$ entry, where darker red indicates more positive values and darker blue indicates more negative values. Diagonal entries remain uncolored, as $\frac{\partial p_{j}}{\partial x_{j}}$ is significantly larger than $\frac{\partial p_{j}}{\partial x_{k}}$ for $j\neq k$. Products are indexed first by firm and then by group assignment. For instance, the first firm ($f=1$) produces nine products: the first two belong to the first group ($g=1$), the next two to the second group ($g=2$), and the remaining five to the third group ($g=3$). In each heatmap, rows and columns are marked to indicate transitions in ownership or product grouping when product $j$ differs from product $j-1$ in firm or group assignment.
The patterns of price effects in the figure align with the economic intuition derived from the simple logit case. First, in all markets, while a firm raises the prices of all its products after improving the attributes of one of its products, the direction of the price effect of other firms’ product attributes depends on industry conduct. [^9] Additionally, products with larger market shares exert stronger effects, as indicated by darker red or blue colors in the figure. That is, under competition ($\phi=0$, panel (a)), prices always decrease in response to improvements in other firms’ product attributes. In contrast, under full profit internalization ($\phi=1$, panel (e)), improvements in the attributes of products owned by colluding partners lead a firm to raise its prices with the magnitude of the price increase being larger for products in the same group.
Second, the degree of profit internalization affects how closely the impact of colluding partners’ product attributes mirrors that of a firm’s own product attributes. When profit internalization is weak ($\phi=0.2$, panel (b)), some product attributes of colluding partners have effects opposite in direction to those of the firm’s own product attributes. Specifically, a firm reduces the prices of its products that are in the same group as (and hence closer substitutes for) those of its colluding partners whose attributes are improved. [^10] Under a random coefficient utility specification where $\nu_{ir}$ in the indirect utility function (1) is heterogeneous, products with similar values of $x_{jt}^{(r)}$ serve as closer substitutes. In this case, when a firm weakly internalizes the profits of its colluding partners, it is more likely to lower the prices of products that share similar characteristics with those of its colluding partners when their attributes improve. For instance, improvements in the attributes of products in the first group ($g=1$), offered by firms 1 and 3, lead to a price reduction for firm 2’s products within the same group, as shown in panel (b), where the corresponding grids are marked in blue. [^11] Figure 1(e) also shows that the price effects are identical for products owned by the same firm and belonging to the same group, as illustrated by the long vertical strips in each heatmap. This uniformity stems from the uniform markup property under the nested logit model specification, as shown in (LABEL:eqn:_foc_nested_logit) in Appendix LABEL:app:_nested_logit; equilibrium markups are identical for products within the same group and offered by the same firm. Accordingly, in the logit case, the price effects are identical for all products owned by the same firm, as shown in Figure LABEL:fig:_heatmap_logit in the Appendix. As $\phi$ increases, the effects of colluding partners’ product attributes on prices of a firm converge to those of the firm’s own product attributes, both in direction and magnitude.
Our key findings – (i) the price effects of own-firm and other-firm attributes are opposite in direction, and (ii) the price effects of other-firm attributes reverse in direction after collusion – are consistently observed across various market structures with different values of $F$, $J$, and $G$, as well as different assignments of products to firms and groups, and varying utility and marginal cost parameters. Moreover, since these findings do not depend on imposing specific choice models, they are expected to remain valid when more flexible models are considered. Building on the analyses discussed thus far, we formalize a testing framework for detecting collusive behavior in the next section.
4 Testing framework